owl shift 7-25-99

Hogan Nguyen (hogann@fnal.gov)
Sun, 25 Jul 1999 06:48:24 -0500

Shift Summary for OWL shift, 25-JULY-99

On Shift: Bob Hsiung, Hogan Nguyen

Hours of beam available at all intensities: 3 pulses ( 15 vac 2pi0's )
Hours of beam available at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 1

Accounting of lost time: 0

========================================================================
Runs ended this shift:

14014 Laser scan

Tapes brought to FCC this shift:

========================================================================
Accesses during shift:

========================================================================
Overview of shift:

Laser scan, cosmic

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14014 1:15 laser scan

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

- Keep taking E832 data: no muon runs, no magnet polarity changes

- The block house humidity is at 9% and the Liebert at 15%. These
cause occasional alarms. Ron Ray has been notified.